Bon wants to take an early retirement in 10 years when he will be twice as old as his daughter Sonia. Sonia is 25 years younger than her father. how old is Bon now? How old will he be when he retires? (use only one variable)(thanks in advance)
Found 2 solutions by ankor@dixie-net.com, MathTherapy:
Answer by ankor@dixie-net.com(22740)

b = Bon's present age
s = Sonia's
:
Write an equation for each statement
Bon wants to take an early retirement in 10 years when he will be twice as old as his daughter Sonia.
b + 10 = 2(s+10)
b + 10 = 2s + 20
b = 2s + 20 - 10
b = 2s + 10
:
Sonia is 25 years younger than her father.
s = b - 25
:
In the 1st simplified equation, replace s with (b-25)
b = 2(b-25) + 10
b = 2b - 50 + 10
b = 2b - 40
40 = 2b - b
b = 40
:
how old is Bon now?
40 yrs old
How old will he be when he retires?
40 + 10 = 50 yrs old
:
;
Check this
Sonia's age: s = 40-25 = 15 yrs: Check in the equation, b = 2s + 10
40 = 2(15) + 10
and
In 10 yrs he will 50, she will 25

Answer by MathTherapy(10551)


Bon wants to take an early retirement in 10 years when he will be twice as old as his daughter Sonia. Sonia is 25 years younger than her father. how old is Bon now? How old will he be when he retires? (use only one variable)(thanks in advance)
The father will be twice his daughter's age when the daughter reaches the age the father was when she was born.
That time is in 10 years, when the daughter will be 25, and the father, 2(25), or 50 years-old.
As the above event will be in 10 years' time, the father is now 50 - 10, or 40 years-old
Since you state that you need only 1 variable used, then we do the following:
Let Bon's age be B
Then his daughter's age is B - 25
We then get: B + 10 = 2(B - 25 + 10)______B + 10 = 2(B - 15)_____B + 10 = 2B - 30
B - 2B = - 30 - 10
- B = - 40
B, or Bon is:
Can you now find out what his age will be in 10 years' time? You should be! I told you before!